OpenStudy (anonymous):

In a finite group z/5z (mod 5), why 2^-1=3 ?

OpenStudy (anonymous):

We say that : \[a^{-1}=b\] iff \[ab=1\] Now we have : \[2\times 3=6=5+1=0+1=1\pmod{5} \] so : \[2^{-1}=3\]

OpenStudy (anonymous):

oh, so basically 2 times what number will give me 1?

OpenStudy (anonymous):

Yeah now i get it. Thanks a bunch

myininaya (myininaya):

I like euclidean's algorithm but what @Noura11 has done is really classy also \[2b=1 \mod 5 \\ 2b-1=5k \\ 2b-5k=1 \\ 5=2(2)+1 \\ 2=1(2) + 0 \\ 5-2(2)=1 \\ 5+2(-2)=1 \\ b=-2 \\ \text{ if you get the inverse is negative } \\ \text{ I always keep adding +5 since we have \mod 5} \\ b=-2+5=3 \]

myininaya (myininaya):

you know until we get the first positive number

OpenStudy (anonymous):

Looks like euclidean's algorithm might come in handy for big numbers.

OpenStudy (anonymous):

@Ankh Yes. Exactly !

OpenStudy (anonymous):

Thank you both!